OpenStudy (anonymous):

help on this word problem Dominique can ice cupcakes twice as fast as Houston. When they work together, Dominique and Houston can ice a large order of cupcakes in 5 hours. How many hours would it take Houston to ice them by himself? 1 4 thirteen-halves fifteen-halves

OpenStudy (tkhunny):

Please write down names for the two missing rates.

OpenStudy (anonymous):

how would i write this in a equation to find the answer??

OpenStudy (anonymous):

5=2x+x ??? im really bad at these

OpenStudy (mertsj):

Dominique takes x hours to do the job Houston takes 2x hours to do the job In 1 hour Dominique gets 1/x of the job done In 1 hour Houston gets 1/(2x) of the job done Since it takes them 5 hours to get the job done, in one hour they do 1/5 of the job. So the part Houston does in 1 hour plus the part Dominique does in 1 hour = 1/5

OpenStudy (mertsj):

\[\frac{1}{x}+\frac{1}{2x}=\frac{1}{5}\]

OpenStudy (jdoe0001):

Dominique takes x hours to do the job <---- wouldn't that give how long Dominique would take if she were to do it by herself?

OpenStudy (mertsj):

yes

OpenStudy (jdoe0001):

seems to me none of the choices match, or there's a typo somewhere

OpenStudy (anonymous):

no thats exactly how it says :/

OpenStudy (jdoe0001):

hmmm

OpenStudy (mertsj):

\[\frac{1}{x}\times10x+\frac{1}{2x}\times10x=\frac{1}{5}\times10x\] 10+5=2x \[15=2x\] \[x=7.5 hours for Dominique to do the job by herself 2x=15 hours for Houston to do the job by himself

OpenStudy (jdoe0001):

yeap... 15 is about right... but is none of the choices, for Houston I had one of these yesterday, and the choices matched the other gal's rate as well

OpenStudy (jdoe0001):

so the choices do fit, Domique's rate, but not Houston's

OpenStudy (mertsj):

I'm guessing some graduate student got the x and 2x mixed up because of the way the problem is worded.

OpenStudy (anonymous):

so the answer is 15/2

OpenStudy (mertsj):

That's the closest answer. That is actually Dominique's time but Houston's is not a choice. So is this a test or what?

OpenStudy (anonymous):

yes for my virtual school. ty

OpenStudy (mertsj):

Well choose 15/2 and see if it is right.

OpenStudy (mertsj):

Was it right?

OpenStudy (anonymous):

i have 2 more question till i've finished the test but i will message u the answer :)

OpenStudy (mertsj):

ok. ty

OpenStudy (anonymous):

yes it was 15/2

OpenStudy (mertsj):

ty